# RouterBench: A Benchmark for Multi-LLM Routing System

### v1

@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable capabilities in addressing a wide range of tasks across academic and industrial scenarios (Bubeck et al., 2023). This has motivated both researchers and practitioners to introduce new LLMs, designed for both generic and specialized use cases, on a near-daily basis  [^1] As of January 16th, 2024, there are 469,848 models listed on huggingface.com. However, the proliferation of LLMs presents a challenge for LLM application builders to identify the most suitable model for their applications. While some proprietary models such as GPT-4 are distinguished by their superior performance, they often incur high economic costs due to the expensive API prices.
@@
Many prior works focus on improving the capabilities of individual LLMs while maintaining low costs. Techniques such as prompting (Wei et al., 2023), quantization (Lin et al., 2023; Kim et al., 2023), and system optimization (Kwon et al., 2023) may reduce a single model’s serving cost, yet with new models emerging daily, these approaches may not remain feasible or scalable in long term. Moreover, the diversity of choices of LLMs available at various price and performance tiers can be daunting for users attempting to select and optimize an appropriate model [^2] As of January 29th, 2024, there are 22,177 language models with 7 billion parameters listed on huggingface.com.
@@
An alternative solution aims to select to optimal LLM for each input through ”routing.” (Chen et al., 2023; Shnitzer et al., 2023; Šakota et al., 2023). Routing offers several advantages over single-LLM optimization. First, it is a lightweight process, which treats each LLM as an input-output black box, avoiding the need to delve into intricate infrastructure details, thus making it flexible and broadly applicable. Second, routing systems benefit from the diversity of LLMs, while single-LLM methods may struggle to keep pace with the expanding LLMs landscape. Lastly, while single-LLM strategies often face a compromise between performance and other factors such as per-token costs, routing systems adeptly balance a spectrum of user demands (Chen et al., 2023; Lee et al., 2023; Lu et al., 2023).
@@
The rise in routing-related research has improved cost-efficiency, enhanced performance, and broadened accessibility. Despite these advances, a comprehensive benchmark for evaluating routing techniques remains absent. We introduce RouterBench, the first comprehensive benchmark designed specifically for assessing router mechanisms in terms of inference dollar cost and performance. RouterBench encompasses a diverse array of tasks and domains, with pre-generated LLM response and quality metrics, on which different routing mechanisms can be efficiently tested without inference. Our experiments revealed that while some previous routing mechanisms have difficulty generalizing to complex tasks and up-to-date models, there are several promising fields on which even simple routing demonstrated outstanding performance.
@@ 2 Related Work
Single LLM Enhancement Fine-tuning is used to improve models for specific tasks, which requires additional training and domain-specific data (Rafailov et al., 2023). Prompting mechanisms like Chain-of-Thought (CoT) (Wei et al., 2023; Zhou et al., 2023; Wang et al., 2023a) and Tree of Thoughts (ToT) (Yao et al., 2023) could bolster LLM performance without additional fine-tuning. Mixture-of-Experts (MoE) (Eigen et al., 2014; Shazeer et al., 2017; Fedus et al., 2022; Du et al., 2022; Shen et al., 2023; Si et al., 2023) is another line of work that explores routing within the model to enhance performance efficiently, which contains specialized ”experts” and routes the input to the best expert. Nevertheless, these single-LLM enhancements are usually model and scenario specific, and could not benefit from the explosion in the number of LLMs.
@@
Routing Unlike LLM Synthesis, routing can select the suitable model for specific input without performing inference on every candidate model. Routing can be classified into two categories, non-predictive routing and predictive routing. Non-predictive routing strategies retrieve outputs from LLMs and directly pick one without a model-assisted synthesis step. Several studies (Madaan et al., 2023; Yue et al., 2023; Lee et al., 2023; Chen et al., 2023) have explored systems that integrate small language models with LLMs. Another methodology involves a layered inference framework, re-routing more complex queries to an advanced model for improved results (Wang et al., 2023b). Predictive routing selects the optimal LLM without requiring to evaluate the output. One line of research has implemented routers utilizing supervised learning algorithms (Shnitzer et al., 2023), while some others use reward model-based techniques (Hari & Thomson, 2023; Lu et al., 2023). Furthermore, meta-model, trained on inputs along with a model-specific token to predict the performance score, represents another approach to determining the most appropriate LLM for use (Šakota et al., 2023). In short, predictive routers could bring substantial cost and performance improvement without sacrificing latency, with a number of early works dedicated to this field.
@@ 3.1 Setup and Basic Operations
The parameters $\theta$ typically include maximum price the user is willing to pay, the desired latency, or number of layers of neural networks for the router model, etc. More details of router parameters will be elaborated and discussed in Section 5.1.
@@
Figure 2: Left: linear interpolation is the process of achieving the cost-performance trade-off between any concrete routers. Point A and B are routers with different input parameters. To achieve the average of A and B, we build router C which routes to A or B with 50% probability each, and it performs the average of A and B in expectation. Middle: Consider points A to E, we can construct the non-decreasing convex hull consists of points A, B, and C. D and E as they can be replaced by strictly superior affine combination of A, B, and C. Right: ABC and DEF are two routing systems (already convexified with ABC extrapolated to (0.1,0) for fair comparison). To compare, we interpolate A and B to $c_{min}=0.1$ and $c_{max}=0.8$, respectively, and then calculate the area under the curve normalized by $c_{max}-c_{min}$ to derive AIQ.
@@
Extrapolation To ensure all routers can enrich our mathematical framework, we also introduce the extrapolation operation, which enables all routers to extend to cost domain $[0,\infty]$. For a given router $R_{\theta}$, we can trivially add more cost to the system without adding performance (for example repeat LLM generation $k$ times and only take the last generation as final output) and thus extend the cost to $\infty$. To extend the router to a smaller cost domain, we simply interpolate the null router (zero cost, zero performance) and $R_{\theta_{1}}$. Thus we are able to achieve any cost level between $[0,\infty]$ to when comparing routers with different domains.
@@ 3.3 Comparing Different Routing Systems
Let’s consider two different routing systems (for example KNN and MLP-based routers), $R_{\theta}$ where $\theta\in{\Theta}$, and $R_{\lambda}$ where $\lambda\in\Lambda$. To compare their effectiveness, we parametrize them by sampling from $\Theta,\Lambda$ to generate a set of points: $R_{\theta_{1}},\dots,R_{\theta_{k}}$, and $R_{\lambda_{1}},\dots,R_{\lambda_{k}}$. Then, we construct non-decreasing convex hull for both groups, $\widetilde{R_{\theta}}$ and $\widetilde{R_{\lambda}}$, defined on a shared domain $[c_{min},c_{max}]$.
We define $AIQ$ (Average Improvement in Quality) for one of the routing system as follows:
@@ 4.2 Benchmark Dataset
RAG Dataset: Additionally, we collected 657 question-ground truth pairs from various major news platforms as well as Wikipedia. To generate the pairs, we ask GPT-4 to come up with questions that can be answered based on the given factual article and the answers, thus generating a dataset to test routers’ performance on retrieval-augmented generation tasks. This initiative is designed to assess the routers’ performance in a complex ”compound system” setting – determining whether routers can adeptly navigate when retrieval abilities are also in play. For instance, when dealing with news published after the GPT-4 knowledge cutoff, it is expected that routers will more frequently opt for models that possess the functionality to access and search the latest internet-based information (e.g., the sonar-medium-online model).
@@ 4.4 A Pilot Study: The Oracle Router
We assessed the performance of various models across the eight datasets, with more details in ( A.5 and 7) while aggregate results are illustrated in Figure 3. The Oracle represents the best possible router: the one that always routes to the best-performing LLM (if there are multiple of them, then route to the cheapest one).
Result: We note that the Oracle router achieves near-optimal performance at a low cost, highlighting the potential for efficient routing among LLMs. Although proprietary models like GPT-4 offer superior performance, their higher cost compared to open-source alternatives is a significant drawback. Other factors such as overalignment could also hurt the generation quality of proprietary models such as Claude 2 (refer to Appendix C). The heatmap in Figure 3 illustrates that, despite WizardLM-13B and Mistral-7B achieving only about 50% accuracy across tasks, their affordability leads to frequent selection by the Oracle, which prioritizes them when they provide correct responses. Moreover, the surprising observation that GPT-4 is seldom chosen suggests the existence of less expensive LLMs that can deliver high-quality answers for most queries. This underscores the substantial opportunity for enhancing LLM systems through cost-effective routing without sacrificing quality.
Figure 3: Left: Accuracy vs Total Cost of all the $11$ LLMs on RouterBench. Right: The Oracle LLMs selection frequency across the $8$ subsets in RouterBench.
@@ 5.1 Predictive Router
We propose a novel set of predictive routers, which do not require pre-generation of LLM outputs. Specifically, we introduce a router $R:x_{i}\to\text{LLM}$, constructed as follows: for an input $x_{i}$, the performance score for $LLM_{j}$ is calculated via:
@@
$P$ denotes the predicted performance of $LLM_{j}$ on sample $x_{i}$, with $\lambda$ representing the willingness to pay (WTP) parameter that delineates the cost-performance trade-off. A higher $\lambda$ indicates a preference for superior performance at a higher cost. We approximate total cost using the cost per token metric. The routing decision for the predictive router is thus formulated as selecting the $LLM$ that optimizes the performance score.
@@
Specifically, the KNN router estimates $\text{performance score}_{ij}$ by identifying the $k$ nearest samples in the training set $D_{train}$ and opting for $LLM_{i}$ demonstrating optimal performance within this subset.
@@
Similarly, for MLP router, we have trained a set of MLP models to predict performance
@@ 5.2 Non-Predictive Routers
This category of routers generates answers from a sequence of Large Language Models (LLMs), evaluates these answers, and bases routing decisions on the evaluation outcomes. Drawing inspiration from (Chen et al., 2023; Wang et al., 2023b), we introduce a cascading router comprising of a total cost parameter $T$, and a sequence of $m$ LLMs, denoted as $LLM_{j}:\text{text}\to\text{text}$, ranked from the least to the most expensive in terms of computational cost and expected accuracy. A key component of its operation is a scoring function $g:\text{text}\to[0,1]$ paired with a threshold $t$ (the ”judge”). Upon receiving a request, it is initially processed by $LLM_{1}$. If $g(o_{1})>t$, the output $o_{1}$ is selected and the process terminates; otherwise, if the cumulative cost is still less than total cost $T$, the router proceeds to the next LLM in the sequence, and returns the current output if not.
Although developing an effective scoring function $g$ for a specific task in a production setting presents challenges, within the context this paper, the router possesses perfect knowledge of the final score, enabling it to consistently select the most cost-effective model that yields a satisfactory response (akin to an oracle). To simulate real-world performance more accurately, we introduce an error parameter $\epsilon\in[0,1]$. The adjusted scoring function $g_{\epsilon}(o)$ is defined as:
@@
A variant of non-predictive router is overgenerate-and-rerank, which generates all potential outcomes from the LLM, assesses each, and outputs the optimal one as determined by a designated reward function. Although its practical application is limited due to significant costs, we will present its results for demonstration.
@@ 5.3 Main Results
Figure 4: Total Cost vs Performance for eleven models and KNN, MLP and Zero routers on RouterBench except MT-Bench. For KNN and MLP, we tested different hyper-parameters and the optimal results are displayed above. The AIQ values are calculated for all $3$ routers. NDCH stands for non-decreasing convex hull, represented by the solid lines. Dotted lines connects points with increasing willingness to pay.
Predictive Router With the KNN and MLP router design, we present the performances of predictive routers across all tasks (other than MT-Bench). The dataset for each task is randomly partitioned into two splits, where the routers are trained on $70\%$ and evaluated on the rest $30\%$. We exclude MT-Bench in this set of experiments due to its limited size to perform such a train-test partition. As shown in Figure 4, both KNN routers and MLP routers achieve the level of performance to the best individual LLMs with lower or similar costs, demonstrating the effectiveness of the proposed routing solutions, despite their simplicity. However, none of the routing algorithms significantly outperform the baseline Zero router (The routers exhibit higher AIQ than the Zero router for MMLU and Winogrande, achieved comparable AIQ for Hellaswag and GSM8K, and underperform on ARC-Challenge and MBPP), the oracle router consistently exceeds all other routers and LLMs in performance, underscoring the room for further advancements in routing algorithms design.
@@
Figure 5: Total Cost vs Performance for eleven models and cascading routers on MMLU, MBPP, and GSM8K. Different error rates are tested, and the AIQ value is computed for Zero Router and zero error rate cascading router. The solid lines represent non-decreasing convex hull and the dotted line represents points with increasing maximum cost parameter.
@@
Cascading Router We present results for cascading routers on MMLU, MBPP, and GSM8K in Figure 5. The results indicate that with each error rate, as the total cost T increases, the cascading router’s performance improves due to the availability of a larger budget for selecting more appropriate models. For lower error rates, the cascading router demonstrates superior performance compared to the Zero router, as evidenced by the higher AIQ value. The router with a zero error rate judge quickly approximates the performance of the Oracle at the same cost and achieves comparable results as the cost further increases. Figure 5 illustrates the cascading routers’ effectiveness, showing they surpass both individual LLMs and the Zero router by a significant margin when the router’s judge has an error rate of up to $0.1$. This indicates the routing technique’s potential when paired with an effective judge.
@@
However, as the judge’s error rates increase, the performance of the cascading router may deteriorate rapidly, particularly when the error rate exceeds $0.2$. Achieving a sufficiently low error rate for certain real-world tasks to benefit from cascading routers might be challenging. Additionally, the sequence in which LLMs are chosen plays a crucial role in performance and offers room for optimization (Chen et al., 2023). Our findings present a simulated upper limit for this method, highlighting the significant potential and the necessity of exploring the optimal implementation of cascading routers for specific applications.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have exhibited remarkable capabilities in addressing a wide range of tasks across academic and industrial scenarios (Bubeck et al., 2023). This has motivated both researchers and practitioners to introduce new LLMs, designed for both generic and specialized use cases, on a near-daily basis  [^1] As of January 16th, 2024, there are 469,848 models listed on huggingface.com. However, the proliferation of LLMs presents a challenge for LLM application builders to identify the most suitable model for their applications. While some proprietary models, such as GPT-4, are distinguished by their superior performance, they often incur high economic costs due to the high API prices.
@@
Many prior works focus on improving the capabilities of individual LLMs while maintaining low costs. Techniques such as prompting (Wei et al., 2022), quantization (Lin et al., 2023; Kim et al., 2023), and system optimization (Kwon et al., 2023) may reduce a single model’s serving cost, yet with new models emerging daily, these approaches may not remain feasible or scalable in long term. Moreover, the diversity of choices of LLMs available at various price and performance tiers can be daunting for users attempting to select and optimize an appropriate model [^2] As of January 29th, 2024, there are 22,177 language models with 7 billion parameters listed on huggingface.com.
@@
An alternative solution aims to select to optimal LLM for each input through ”routing.” (Chen et al., 2023; Shnitzer et al., 2023; Šakota et al., 2023). Routing offers several advantages over single-LLM optimization. First, it is a lightweight process, which treats each LLM as an input-output black box, avoiding the need to delve into intricate infrastructure details, thus making it flexible and broadly applicable. Second, routing systems benefit from the diversity of LLMs, while single-LLM methods may struggle to keep pace with the expanding LLM landscape. Lastly, while single-LLM strategies often face a compromise between performance and other factors such as per-token costs, routing systems adeptly balance a spectrum of user demands.
@@
The rise in routing-related research has improved cost efficiency, enhanced performance, and broadened accessibility (Chen et al., 2023; Lee et al., 2023; Lu et al., 2023). Despite these advances, a comprehensive benchmark for evaluating routing techniques remains absent. We introduce RouterBench, the first comprehensive benchmark designed specifically for assessing router mechanisms in terms of inference dollar cost and performance. RouterBench encompasses a diverse array of tasks and domains, with pre-generated LLM response and quality metrics, on which different routing mechanisms can be efficiently tested without inference. Our experiments revealed that while some previous routing mechanisms have difficulty generalizing to complex tasks and up-to-date models, there are several promising fields on which even simple routing demonstrated outstanding performance.
@@ 2 Related Work
Single LLM Enhancement Fine-tuning is used to improve models for specific tasks, which requires additional training and domain-specific data (Rafailov et al., 2023). Prompting mechanisms like Chain-of-Thought (CoT) (Wei et al., 2022; Zhou et al., 2023; Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) could bolster LLM performance without additional fine-tuning. Mixture-of-Experts (MoE) (Eigen et al., 2014; Shazeer et al., 2017; Fedus et al., 2022; Du et al., 2022; Shen et al., 2023; Si et al., 2023) is another line of work that explores routing within the model to enhance performance efficiently, which contains specialized ”experts” and routes the input to the best expert. Nevertheless, these single-LLM enhancements are usually model and scenario specific, and could not benefit from the explosion in the number of LLMs.
@@
Routing Unlike LLM Synthesis, routing can select the suitable model for specific input without performing inference on every candidate model. Routing can be classified into two categories, non-predictive routing and predictive routing. Non-predictive routing strategies retrieve outputs from LLMs and directly pick one without a model-assisted synthesis step. FrugalGPT (Chen et al., 2023) presents an inaugural application of this type of strategy, which employs a generation judger that assesses the quality of responses from various LLMs to a given query, invoking LLMs sequentially until an answer meets a predefined quality threshold. Several studies (Madaan et al., 2023; Yue et al., 2023; Lee et al., 2023) also have explored systems that integrate small language models with LLMs. Another methodology involves a layered inference framework, re-routing more complex queries to an advanced model for improved results (Wang et al., 2023). Predictive routing selects the optimal LLM without requiring to evaluate the output. One line of research has implemented routers utilizing supervised learning algorithms (Shnitzer et al., 2023), while some others use reward model-based techniques (Hari & Thomson, 2023; Lu et al., 2023). Furthermore, meta-model, trained on inputs along with a model-specific token to predict the performance score, represents another approach to determining the most appropriate LLM for use (Šakota et al., 2023). In short, predictive routers could bring substantial cost and performance improvement without sacrificing latency, with a number of early works dedicated to this field.
@@ 3.1 Setup and Basic Operations
The parameters $\theta$ typically include the maximum price the user is willing to pay, the desired latency, or a number of layers of neural networks for the router model, etc. More details of router parameters will be elaborated and discussed in Section 5.1.
@@
Figure 2: Left: linear interpolation is the process of achieving the cost-performance trade-off between any concrete routers. Point A and B are routers with different input parameters. To achieve the average of A and B, we build router C which routes to A or B with 50% probability each, and it performs the average of A and B in expectation. Middle: Consider points A to E, we can construct the non-decreasing convex hull consisting of points A, B, and C. D and E as they can be replaced by a strictly superior affine combination of A, B, and C. Right: ABC and DEF are two routing systems (already convexified with ABC extrapolated to (0.1,0) for a fair comparison). To compare, we interpolate A and B to $c_{min}=0.1$ and $c_{max}=0.8$, respectively, and then calculate the area under the curve normalized by $c_{max}-c_{min}$ to derive AIQ.
@@
Extrapolation To ensure all routers can enrich our mathematical framework, we also introduce the extrapolation operation, which enables all routers to extend to the cost domain $[0,\infty]$. For a given router $R_{\theta}$, we can trivially add more cost to the system without adding performance (for example repeat LLM generation $k$ times and only take the last generation as final output) and thus extend the cost to $\infty$. To extend the router to a smaller cost domain, we simply interpolate the null router (zero cost, zero performance) and $R_{\theta_{1}}$. Thus we are able to achieve any cost level between $[0,\infty]$ to when comparing routers with different domains.
@@ 3.3 Comparing Different Routing Systems
Let’s consider two different routing systems (for example KNN and MLP-based routers), $R_{\theta}$ where $\theta\in{\Theta}$, and $R_{\lambda}$ where $\lambda\in\Lambda$. To compare their effectiveness, we parametrize them by sampling from $\Theta,\Lambda$ to generate a set of points: $R_{\theta_{1}},\dots,R_{\theta_{k}}$, and $R_{\lambda_{1}},\dots,R_{\lambda_{k}}$. Then, we construct a non-decreasing convex hull for both groups, $\widetilde{R_{\theta}}$ and $\widetilde{R_{\lambda}}$, defined on a shared domain $[c_{min},c_{max}]$.
We define $AIQ$ (Average Improvement in Quality) for one of the routing systems as follows:
@@ 4.2 Benchmark Dataset
RAG Dataset: To evaluate routers in a more practical setting, we collected 800 user queries from one of Martian’s clients, an LLM-assisted search company, and constructed an RAG dataset based on these queries. These queries cover topics including sports, history, media & art, and politics, and all have ground truth answers. We then manually collected the ground truths, which were used to evaluate answers from selected groups of LLM and LLM-assisted search engines. This initiative is designed to assess the routers’ performance in a complex ”compound system” setting (Zaharia et al., 2024) – determining whether routers can adeptly navigate when retrieval abilities are also in play. For instance, when dealing with news published after the GPT-4 knowledge cutoff, routers are expected to more frequently opt for models that can access and search the latest internet-based information (e.g. sonar-medium-online).
@@ 4.3 Dataset Construction Process
Figure 3: Left: Accuracy vs Total Cost of all the $11$ LLMs on RouterBench. Right: The Oracle LLMs selection frequency across the $7$ subsets in RouterBench.
@@ 4.4 A Pilot Study: The Oracle Router
We assessed the performance of various models across the eight datasets, with more details in ( A.4 and 7) while aggregate results are illustrated in Figure 3. The Oracle represents the best possible router: the one that always routes to the best-performing LLM (if there are multiple of them, then route to the cheapest one).
Result: We note that the Oracle router achieves near-optimal performance at a low cost, highlighting the potential for efficient routing among LLMs. Although proprietary models like GPT-4 offer superior performance, their higher cost than open-source alternatives is a significant drawback. Factors such as overalignment could also hurt the generation quality of proprietary models such as Claude 2 (refer to Appendix C). The heatmap in Figure 3 illustrates that, despite WizardLM-13B and Mistral-7B achieving only about 50% accuracy across tasks, their affordability leads to frequent selection by the Oracle, prioritizing them when they provide correct responses. Moreover, the surprising observation that GPT-4 is seldom chosen suggests the existence of less expensive LLMs that can deliver high-quality answers for most queries. This underscores the substantial opportunity for enhancing LLM systems through cost-effective routing without sacrificing quality.
@@ 5.1 Predictive Router
We propose a novel set of predictive routers that do not require the pre-generation of LLM outputs. Specifically, we introduce a router $R:x_{i}\to\text{LLM}$, constructed as follows: for an input $x_{i}$, the performance score for $LLM_{j}$ is calculated via:
@@
$P$ denotes the predicted performance of $LLM_{j}$ on sample $x_{i}$, with $\lambda$ representing the willingness to pay (WTP) parameter that delineates the cost-performance trade-off. A higher $\lambda$ indicates a preference for superior performance at a higher cost. We approximate the total cost using the cost per token metric. The routing decision for the predictive router is thus formulated as selecting the $LLM$ that optimizes the performance score.
@@
Specifically, the KNN router estimates $\text{performance score}_{ij}$ by identifying the $k$ nearest samples in the training set $D_{train}$ and opting for $LLM_{i}$, demonstrating optimal performance within this subset.
@@
Figure 4: Total Cost vs. Performance for eleven models and KNN, MLP, and Zero routers on RouterBenchexcept for MT-Bench. For KNN and MLP, we tested different hyper-parameters, and the optimal results are displayed above. The AIQ values are calculated for all $3$ routers. NDCH stands for non-decreasing convex hull, represented by the solid lines. Dotted lines connect points with increasing willingness to pay.
Similarly, for the MLP router, we have trained a set of MLP models to predict the performance
@@ 5.2 Non-Predictive Routers
This category of routers generates answers from a sequence of Large Language Models (LLMs), evaluates these answers, and bases routing decisions on the evaluation outcomes. Drawing inspiration from (Chen et al., 2023; Wang et al., 2023), we introduce a cascading router comprising of a total cost parameter $T$, and a sequence of $m$ LLMs, denoted as $LLM_{j}:\text{text}\to\text{text}$, ranked from the least to the most expensive in terms of computational cost and expected accuracy. A key component of its operation is a scoring function $g:\text{text}\to[0,1]$ paired with a threshold $t$ (the ”judge”). Upon receiving a request, it is initially processed by $LLM_{1}$. If g(o1)¿t, the output o1 is selected, and the process terminates; otherwise, if the cumulative cost is still less than the total cost T, the router proceeds to the next LLM in the sequence and returns the current output if not.
Although developing an effective scoring function $g$ for a specific task in a production setting presents challenges, within the context of this paper, the router possesses perfect knowledge of the final score, enabling it to consistently select the most cost-effective model that yields a satisfactory response (akin to an oracle). To simulate real-world performance more accurately, we introduce an error parameter $\epsilon\in[0,1]$. The adjusted scoring function $g_{\epsilon}(o)$ is defined as:
@@
A variant of the non-predictive router is overgenerate-and-rerank, which generates all potential outcomes from the LLM, assesses each, and outputs the optimal one as determined by a designated reward function. Although its practical application is limited due to significant costs, we will present its results for demonstration.
@@ 5.3 Main Results
Figure 5: Total Cost vs Performance for eleven models and cascading routers on MMLU, MBPP, and GSM8K. Different error rates are tested, and the AIQ value is computed for Zero Router and zero error rate cascading router. The solid lines represent the non-decreasing convex hull and the dotted line represents points with increasing the maximum cost parameter.
Predictive Router With the KNN and MLP router design, we present the performances of predictive routers across all tasks (other than MT-Bench). The dataset for each task is randomly partitioned into two splits, where the routers are trained on $70\%$ and evaluated on the rest $30\%$. We exclude MT-Bench in this set of experiments due to its limited size in performing such a train-test partition. As shown in Figure 4, both KNN routers and MLP routers achieve the level of performance to the best individual LLMs with lower or similar costs, demonstrating the effectiveness of the proposed routing solutions, despite their simplicity. However, none of the routing algorithms significantly outperform the baseline Zero router (The routers exhibit higher AIQ than the Zero router for MMLU and Winogrande, achieved comparable AIQ for Hellaswag and GSM8K, and underperform on ARC-Challenge and MBPP), the oracle router consistently exceeds all other routers and LLMs in performance, underscoring the room for further advancements in routing algorithms design.
@@
Cascading Router We present results for cascading routers on MMLU, MBPP, and GSM8K in Figure 5. The results indicate that with each error rate, as the total cost T increases, the cascading router’s performance improves due to the availability of a larger budget for selecting more appropriate models. For lower error rates, the cascading router performs better than the Zero router, as evidenced by the higher AIQ value. The router with a zero error rate judge quickly approximates the performance of the Oracle at the same cost and achieves comparable results as the cost further increases. Figure 5 illustrates the cascading routers’ effectiveness, showing they surpass both individual LLMs and the Zero router by a significant margin when the router’s judge has an error rate of up to $0.1$. This indicates the routing technique’s potential when paired with an effective judge.
@@
However, as the judge’s error rates increase, the performance of the cascading router may deteriorate rapidly, particularly when the error rate exceeds $0.2$. Achieving a sufficiently low error rate for certain real-world tasks to benefit from cascading routers might be challenging. Additionally, the sequence in which LLMs are chosen plays a crucial role in performance and offers room for optimization (Chen et al., 2023). Our findings present a simulated upper limit for this method, highlighting the potential and the necessity of exploring the optimal implementation of cascading routers for specific applications.